\begin{thebibliography}{5}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[{\citenamefont{{C. Zhang, V. W. Scarola, S. Tewari, and S. Das
  Sarma}}(2007)}]{Zhang07}
\bibinfo{author}{\bibnamefont{{C. Zhang, V. W. Scarola, S. Tewari, and S. Das
  Sarma}}}, \bibinfo{journal}{Proc. Natl. Acad. Sci. U.S.A.}
  \textbf{\bibinfo{volume}{104}}, \bibinfo{pages}{18415}
  (\bibinfo{year}{2007}).

\bibitem[{\citenamefont{Kitaev}(2006)}]{Kitaev06}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Kitaev}},
  \bibinfo{journal}{Ann. Phys.} \textbf{\bibinfo{volume}{321}},
  \bibinfo{pages}{2} (\bibinfo{year}{2006}).

\bibitem[{\citenamefont{Schmidt et~al.}()\citenamefont{Schmidt, Dusuel, and
  Vidal}}]{Schmidt08}
\bibinfo{author}{\bibfnamefont{K.~P.} \bibnamefont{Schmidt}},
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Dusuel}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Vidal}},
  \bibinfo{note}{arXiv:0709.3017 (accepted for publication in Phys. Rev.
  Lett.)}.

\bibitem[{\citenamefont{Baskaran et~al.}(2007)\citenamefont{Baskaran, Mandal,
  and Shankar}}]{Baskaran07}
\bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{Baskaran}},
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Mandal}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Shankar}},
  \bibinfo{journal}{Phys. Rev. Lett.} \textbf{\bibinfo{volume}{98}},
  \bibinfo{pages}{247201} (\bibinfo{year}{2007}).

\bibitem[{\citenamefont{Dusuel et~al.}()\citenamefont{Dusuel, Schmidt, and
  Vidal}}]{Dusuel08}
\bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Dusuel}},
  \bibinfo{author}{\bibfnamefont{K.~P.} \bibnamefont{Schmidt}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Vidal}},
  \bibinfo{note}{in preparation}.

\end{thebibliography}

\documentclass[prl,twocolumn,superscriptaddress,floatfix]{revtex4}
\usepackage{graphicx,amsfonts,amssymb,amsmath,hyperref}

\newif\ifhyper
\hypertrue
\ifhyper
\hypersetup{
   citecolor = {green},
   colorlinks = {true}, 
   urlcolor = {blue} 
}
\fi

\begin{document}

\title{Comment on ``Anyonic braiding in optical lattices"}

\author{Julien Vidal}
\email{vidal@lptmc.jussieu.fr}
\affiliation{Laboratoire de Physique Th\'eorique de la Mati\`ere Condens\'ee,
  CNRS UMR 7600, Universit\'e Pierre et Marie Curie, 4 Place Jussieu, 75252
  Paris Cedex 05, France}

\author{S\'ebastien Dusuel}
\email{sdusuel@gmail.com}
\affiliation{Lyc\'ee Louis Thuillier, 70 Boulevard de Saint Quentin,
  80098 Amiens Cedex 3, France}

\author{Kai Phillip Schmidt}
\email{schmidt@fkt.physik.uni-dortmund.de}
\affiliation{Lehrstuhl f\"ur theoretische Physik, Otto-Hahn-Stra\ss e 4, D-44221
Dortmund, Germany}

\maketitle

In Ref.~\cite{Zhang07}, C. Zhang {\it et al.} proposed an experimental scheme to detect the braiding statistics in Kitaev's honeycomb model \cite{Kitaev06}. 
Although the perspective to observe anyons experimentally is very challenging, we would like to point out some major technical and conceptual mistakes which invalidate the conclusion drawn in \cite{Zhang07}. 

Indeed, the study presented in \cite{Zhang07} relies on a perturbation theory of the Kitaev model in the gapped phase ($J_z \gg J_x,J_y$) \cite{Kitaev06,Schmidt08}. At order 0, the low-energy eigenstates are built from collinear dimer configuration on each $z$-dimer. However, contrary to what is assumed in \cite{Zhang07}, this is no longer true when switching the perturbation on, 
as can be seen for instance in a first order calculation of
the ground state correlation function $\langle \sigma^x_i \sigma^x_j \rangle=\frac{J_x}{2J_z}$ on a $x-$dimer $(i,j)$.
This basic error has led the authors to propose a form of the ground state which gives 
$\langle \sigma^x_i \sigma^x_j \rangle\neq 0$ on a $z$-dimer $(i,j)$ whereas exact results show that all two-spin correlations vanish except
$\langle \sigma^\alpha_i \sigma^\alpha_j \rangle$ on a $\alpha$-dimer \cite{Baskaran07}.

In addition, all spin operations used in \cite{Zhang07} to create anyons do also produce fermionic excitations that are likely to spoil the proposed detection set-up \cite{Dusuel08}. Once again, we emphasize that this effect already occurs at first order in perturbation. 

To summarize, results given in \cite{Zhang07} are only valid for $J_x=J_y=0$ where the model is trivial and all low-energy levels exactly degenerate. In other words, the proposal of measuring anyonic excitations suggested in \cite{Zhang07} is strongly questionable.


\end{document}